OpenStudy (anonymous):

Express the complex number in trigonometric form. 6 - 6i

OpenStudy (anonymous):

6sqrt(2)[ cos(-pi/4)+ i sin(-pi/4)]

OpenStudy (anonymous):

Z=6-6i , the |z| = sqrt(6^2 + 6^2) = 6sqrt(2) Amp(z)=tan^-1(-6/6) = tan^-1(-1) = (-pi/4)

OpenStudy (anonymous):

thats not an answer choice i have 6sqrt(2){cos(5pi/4)+i sin(5pi/4) or 6sqrt(2){cos(7pi/4)+i sin(7pi/4

OpenStudy (anonymous):

@siddarth95

hartnn (hartnn):

all trigonometric ratios are periodic with 2pi so, cos (-pi/4) = cos (-pi/4 + 2pi)=cos (...?...)

OpenStudy (anonymous):

What @hartnn said... Cos(2pi-pi/4)

OpenStudy (anonymous):

i started off by doing [\sqrt{6^2+(-6)^2} and got \sqrt{72} which is 6\sqrt{2} \] then to fid theta i did arctan of (-6/6) which is -45 degrees and that is in the fourth quadrant of the unit circle but now i dont know what to do im stuck @hartnn and my choices are 6sqrt(2){cos(5pi/4)+i sin(5pi/4) or 6sqrt(2){cos(7pi/4)+i sin(7pi/4

hartnn (hartnn):

what is 2pi- pi/4 =... ?